Reconciling Changes in Balance Sheet Accounts

The following table presents selected items from the 2014 and 2013 balance sheets and 2014 income statement of Walgreen Company.

WALGREEN CO. AND SUBSIDIARIES ($ millions)
Selected Balance Sheet Data Selected Income Statement Data
2014 2013 2014
Inventories $16,088 $14,756 Cost of Merchandise Sold $103,384
Property and Equipment, Less Depreciation Expense 1,618
Accumulated Depreciation 23,052 22,368
Trade Accounts Payable 9,620 9,170 Net Earnings 5,428
Retained Earnings 37,754 33,696
a. Compute the cash paid for merchandise inventories in 2014. Assume that trade accounts payable is only for merchandise purchases.
$Answer million
b. Compute the net cost of property acquired in 2014.
$Answer million
c. Compute the cash dividends paid in 2014.
$Answer million
Answer #1

Answers:

Compute the cash paid for merchandise inventories in 2014. Assume that trade accounts a. payable is only for merchandise purc

Working:

(all amounts are in million)

Answer (a):

Purchases = Ending inventory + Cost of goods sold - Beginning inventory = 16088 + 103384 - 14756 = $104,716

Cash paid for merchandise inventories in 2014 = Purchase + Beginning Trade Accounts - Ending Trade Accounts Payable

= 104716 + 9170 - 9620

= $104,266 million

Answer (b):

Net cost of property acquired in 2014 = Ending balance of Net Property and Equipment - (Beginning balance of Net Property and Equipment - Depreciation expense of 2014

= 23,052 - (22,368 - 1,618)

= $2,302 million

Answer (c):

cash dividends paid in 2014 = Beginning balance of Retained earnings + Net earnings - Ending balance of Retained earnings

= 33696 + 5428 - 37754

=$1,370 million